# Chaos controlled and disorder driven phase transitions by breaking permutation symmetry

### v1

@@ Abstract
Figure 2: (a) $\left\langle J^{2}\right\rangle$ with respect to time or number of kicks $n$ for different disorder strengths $w$ with $k=1$, $N=12$ qubits, averaged over 100 disorder realizations. Inset shows the entanglement $S_{N/2}$ as a function of $n$ for the same parameters as in the main figure.
@@
Figure 3: (a) Change in scaling of $\langle J^{2}\rangle$ from $N^{2}/4+N/2$ to $43N/$ 4 as $w$ is increased where the chaos parameter $k$ is set to 1.(b) Same for $S_{N/2}(A)$ as a function of $N$ at $k=1$ as discussed in the text. The red dashed line in Fig. (a) corresponds to $N^{2}/4+N/2$ whereas the black dotted line corresponds to the RMT result $43N/$ 4. Similarly, red dashed line in Fig. (b) corresponds to fitted line $0.41+0.54\log_{2}(N/2+1)$ and the black dotted line is the RMT expression given by $N/2-1/2\ln 2$.
It is tempting to associate this change in the scaling of $\langle J^{2}\rangle$ and $S_{N/2}$ with a second order phase transition. Therefore, we now investigate the possible existence of a critical $w_{c}$ below which the dynamics is predominantly within the permutation symmetric subspace and above which the dynamics captures the properties of the full Hilbert space described by RMT, and obtain relevant critical exponents of the associated phase transition, if any. We therefore propose a finite size scaling of the form Botet and Jullien (1983); Skinner et al. (2019); Lee et al. (2017); Li et al. (2018); Cardy (1988)
${\langle J^{2}\rangle}=N^{\zeta/\nu}F((w-w_{c})N^{1/\nu}),$ (7)
where $\nu$ is the correlation length exponent. Fig. 4 (main) shows the crossing of the data around $w_{c}=2.4$ when $k=1$ whereas the collapse of the data with $\nu=0.59$ and $\zeta=0.59$ is shown in the inset of Fig. 4. In order to confirm this phase transition further, we repeat the calculations for other values of $k$, the results of which are shown in Figure 5. It is to be noted that $k=4$ is a mixed phase space with considerable chaos already present in the system.
@@
Figure 4: $\langle J^{2}\rangle/N^{\zeta/\nu}$ as a function of disorder strength $w$ for $k=1$ and different system sizes. These curves cross each other at $w=w_{c}\approx 2.4$. The inset shows the collapse of the data with $\zeta=0.59$ and $\nu=0.59$ consistent with the scaling proposed in Eq. 7.
@@
Figure 5: (a) The main figure shows the data corresponding to different system sizes for $k=2$ crossing each other at $w=w_{c}\sim 0.41$ whereas the inset shows the collapse of the data with $\nu=0.40$ and $\zeta=0.49$.(b) Same as (a) but for $k=4$ with $w=w_{c}\sim 0.17$ and $\nu=0.3$.
@@
Table 1 lists the critical disorder strength $w_{c}$ along with the critical exponent $\nu$ for different values of the interaction strength or chaos parameter $k$. Intuitively, the chaotic dynamics at larger $k$ values will require smaller disorder to take the dynamics to the full Hilbert space, which is also what we observe. Interestingly, we also find a weak parameter dependence on the critical exponent $\nu$ for this phase transition. The exponent $\nu$ lies within a range of 0.6 to 0.3 for the parameters studied. Such a variation seems to be a characteristic of a disordered system such as the well studied many-body-localization one Wang et al. (2021). We also find that in most cases, $\nu\sim\zeta$. The quality of the data collapse for different values of $k$ confirm the presence of a second order phase transition. At the same time, we do observe that identifying correct $w_{c}$ values for larger $k$ is difficult since $w_{c}$ then is very small and a single crossing point for all system sizes is difficult to obtain. On the other hand when $k\rightarrow 0$, the dynamics can get highly non-ergodic and these parameters need more careful examinations. We have further confirmed this phase transition using the entanglement entropy $S_{N/2}(A)$ with a similar finite size scaling as in Eq. 7 giving $\nu=0.6$, $w_{c}=1.8$ when $k=1$, as shown in Fig. 6.
It is also observed that the collapse is relatively good when the clean system is either predominantly chaotic or predominantly regular. For a mixed phase space, with significant fraction of both regular and chaotic regions, say at $k=3$, the quality of the scaling seems to depend on the initial state. The value in the table 1 for $k=3$ and $k=4$ corresponds to chaotic regions. For regular regions in these cases, we could not observe scaling behaviour. This may be attributed to the small values of angular quantum number $j$, that we are forced to restrict ourselves to, as well as the complexity of mixed phase space (see for example Madhok (2015) for cases when the initial state in a kicked top is at the border between chaos and regular regions) and the influence of phenomena such as chaos assisted tunneling Tomsovic and Ullmo (1994); S.Keshavamurthy and P.Schlagheck ; Krithika et al. (2023), that warrants a detailed separate study.
@@
Table 1: Critical disorder strength and exponents $\nu$, $\zeta$ for different chaos parameter, $k$. We have used $\langle J^{2}\rangle$ with $N=12,14$ and $16$ to identify the critical points with time averaging ranging from $n=10^{5}$ to $3\times 10^{5}$ for 100 different realizations.
@@
Figure 6: (a) Plot of $\langle S_{N/2}(A)\rangle/N^{\zeta/\nu}$ vs $w$ for $k=1$ corresponding to different system sizes cross each other at $w=w_{c}\sim 1.8$. Inset shows the collapse of the data with $\nu=0.67$ and $\zeta=0.76$.
@@
In summary, this work has shown the existence of a critical strength of permutation symmetry breaking, beyond which there is a second order phase transition from area law entangled states with total angular momentum scaling as $N$, to volume law entangled states with total angular momentum scaling as $\sqrt{N}$. The critical disorder strength shows clear dependence on the dynamics being regular or chaotic in the disorder-free permutation symmetric space. The exponents on the other hand seem to show a weaker dependence. Results not shown indicate that the criticality is robust on changing the the initial state except when it is in the regular region of the mixed phase space.
@@
From an applications perspective, our results show that for disorder less than the critical strength, permutation symmetric systems are likely to be reliably used for their intended application, being at least not volume law entangled in the full Hilbert space. It is possible that the proposed phase transition be experimentally observed by extending few qubit experimental kicked top studies Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019), wherein disorder is anyway present. In fact, a disordered all-to-all interacting system, simulating the SYK model, has been recently experimentally studied using a combination of atomic cloud in a cavity and controllable light shift in the large detuning limit Sauerwein et al. (2023).
@@
Among future directions, it is interesting to explore if breaking permutation symmetry through other methods, say by introducing a power law interaction such as in Russomanno et al. (2021), will also exhibit a similar phase transition as induced by disorder.
@@
Acknowledgements: MC and UD acknowledege the HPC facility Chandra at IIT Palakkad where the computations were carried out. AL thanks Sylvia Pappalardi for discussions at an early stage of this work, at the program on “Dynamical Foundations of Many-Body Chaos”, Institut Pascal, Université Paris-Saclay, during March-April 2023.

### v2

@@ III Results and Discussions
Figure 4: Top: $\left\langle J^{2}\right\rangle$ as a function of number of kicks $n$ for different disorder strengths $w$ with $k=1$, $N=12$ qubits, averaged over 100 disorder realizations. Bottom: shows the von Neumann entanglement entropy $S_{N/2}$ as a function of $n$ for the same parameters. The two vertical lines correspond to the two Heisenberg time scales discussed in the text, $t_{PSS}^{H}$ (blue) and $t_{FHS}^{H}$ (green). The black dashed horizontal line in both figures correspond to the RMT values in full Hilbert space.
@@
Figure 5: Top: Change in scaling of $\overline{\langle J^{2}\rangle}$ from $N^{2}/4+N/2$ (red dashed line) to $43N/$ 4 (black dotted line) as $w$ is increased with $k$ set to 1. Bottom: Same for $\overline{S}_{N/2}$ as a function of $N$ at $k=1$ where the red dashed line corresponds to fitted line $0.41+0.54\log_{2}(N/2+1)$ and the black dotted line is the RMT expression given by $N/2-1/2\ln 2$.
Clearly, for a generic initial coherent state, $\langle J^{2}(t)\rangle$ decreases exponentially with time and saturates to a disorder independent, but initial state dependent value, which is $O(N^{2})$. The irreversible nature of the dynamics arises from the disorder averaging even in finite systems. When $\theta=\pi/2$ and $\phi=0$ or $\pi$, there is no decay and $\langle J^{2}(t)\rangle=\langle J^{2}(0)\rangle=j(j+1)$, where $j=N/2$. These special initial states have no evolution except for a phase as they are eigenstates of the Hamiltonian. Interestingly, when $\phi=\pm\pi/2$ (any value of $\theta$), $\langle J^{2}(n)\rangle$ saturates to the RMT value given by $43N/$ 4 for any non-zero value of disorder. These initial states are in the $y-z$ plane of the Bloch sphere, and are maximally coherent (in the sense of quantum information theory, as discussed below) in the preferred $x$ basis as this is the direction of the interaction. These are the only initial states for which the saturation value of $J^{2}$ scales linearly as $N$, rather than quadratically, and are unstable in the sense that any deviation in the $\phi$ value changes this from the RMT scaling. We will find that in the presence of the transverse field kicks when $p\neq 0$, these behaviours are interestingly modified, while some persist. Below, we present our results in the presence of kicks, first for a generic coherent state and then for a non-generic, initial coherent states.
III.1 A generic initial state
We present the results of numerical calculations which are performed with an initial state |$\ket{\psi_{0}}$ in the PSS as given in Eq. 5 with $\theta=2.25,~{}\phi=1.1$ ($\blacktriangle$ in Fig. 2) and is evolved to |$\ket{\psi_{n}}=U_{w}^{n}\ket{\psi_{0}}$ using the Floquet operator $U_{w}$ given in Eq. 2. This initial state is “generic” in the sense that it has no special dynamical significance in the permutation symmetric classical limit. It is still a coherent state and not generic in the sense of Haar measure. We use the fast Walsh Hadamard transform to reach system sizes of $N=16$. Fig. 3 compares the dynamics of the system, both, in the absence and in the presence of kicks for two different values of $w$. We find that in the absence of kicks, and for the generic initial coherent state when $\cos\phi\neq 0$, the system reaches its saturation value when time $t>t^{*}\sim\frac{\sqrt{N}}{wk}$ in the large $N$ limit, which is much smaller than that in the presence of kicks. The relevant time scales in the presence of disorder and kicks is more clearly shown in Fig. 4 using $\langle J^{2}(n)\rangle$ and $S_{N/2}(A)$ for $k=1$. In contrast to the unkicked model, there is a disorder dependent steady state value eventually reached for all values of the disorder; however we find an interesting feature consisting of two regimes, first is till time $n\sim t^{H}_{PSS}=N+1$ corresponding to the Heisenberg time relevant to the PSS, when there is a quasi steady state value. The other is at $n\sim t^{H}_{FHS}=2^{N}$, the Heisenberg time relevant to the FHS where $\langle J^{2}\rangle$ and $S_{N/2}(A)$ again starts saturating, which implies that the system has “realized” the presence of the full Hilbert space. Notice that with large disorder strength such as $w=5$, when the full Hilbert space is being accessed, the time scale at $t^{H}_{FHS}$ has become irrelevant for this quantity. Instead the Heisenberg time for the PSS provides a time scale for thermalization, when $\langle J^{2}(n)\rangle$ and $S_{N/2}(A)$ saturates close to the RMT value corresponding to FHS. This is to be expected as the Heisenberg time for the PSS is also the Ehrenfest or log-time for the dynamics in the FHS. For intermediate values of the disorder strength $w$, the nonequilibrium state’s $J^{2}$ as well as the entanglement saturates to non-thermal (non-RMT) values, signifying some kind of localization. It would be interesting to compare these with the well-studied case of MBL Pal and Huse (2010); Oganesyan and Huse (2007); Luitz et al. (2015).
To quantify the saturation value as a function of the disorder and system size, we now study the behavior of long time averaged values (averaged from $n=10^{5}$ to $n=3\times 10^{5}$) of $J^{2}$ and $S_{N/2}(A)$, denoted as $\overline{\langle J^{2}\rangle}$ and $\overline{S}_{N/2}$, respectively. From the discussions above, we expect that as $w\to 0$ or small when the dynamics is predominantly within the permutation symmetric subspace, the scaling of long time limit of various quantities will follow the PSS scaling, $i.e.,$ $\overline{\langle J^{2}\rangle}\sim N^{2}$ and $\overline{S}_{N/2}\sim\log_{2}(N/2+1)$. In the other extreme limit, when $w\to\infty$ and the dynamics is described by RMT in the full $2^{N}-$dimensions, $\overline{\langle J^{2}\rangle}=3N/4$ and $\overline{S}_{N/2}\sim N/2$. Fig. 5 highlights the above discussed change in scaling when $w$ is increased. Note that for the solvable case in the absence of kicks, while Eq. 9, indicates an earlier saturation time scale $\sim\sqrt{N}$ for generic initial coherent states, the saturation value itself is still similar to the PSS case with $\langle J^{2}\rangle\sim N^{2}$.
@@ III.1 A generic initial state
Figure 6: (Top): $\overline{\langle J^{2}\rangle}/N^{\zeta/\nu}$ as a function of disorder strength $w$ at $k=1$ for different system sizes. These curves cross each other at $w=w_{c}\approx 2.38$. (Bottom): The collapse of the data with $\zeta=0.59$ and $\nu=0.60$ consistent with the scaling proposed in Eq. 10.
@@
Figure 7: (Top): The main figure shows the data corresponding to different system sizes for $k=2$ crossing each other at $w=w_{c}\sim 0.41$ whereas the inset shows the collapse of the data with $\nu=0.40$ and $\zeta=0.49$. (Bottom): Same as above but for $k=4$ with $w=w_{c}\sim 0.17$, $\nu=0.3$ and $\zeta=0.32$.
It is tempting to associate this shift in the scaling of $\overline{\langle J^{2}\rangle}$ and $\overline{S}_{N/2}$ with a second order phase transition. Therefore, we now investigate the possible existence of a critical $w_{c}$ below which the dynamics is predominantly within the PSS and above which the dynamics captures the properties of the FHS described by RMT, and obtain relevant critical exponents of the associated phase transition, if any. We propose a finite size scaling of the form Botet et al. (1982); Botet and Jullien (1983); Skinner et al. (2019); Lee et al. (2017); Li et al. (2018); Cardy (1988)
${\overline{\langle J^{2}\rangle}}=N^{\zeta/\nu}F((w-w_{c})N^{1/\nu}),$ (10)
where $\nu$ is the correlation length exponent. Fig. 6 (top) shows the crossing of the data around $w_{c}=2.38$ when $k=1$ whereas the collapse of the data with $\nu=0.60$ and $\zeta=0.59$ is shown in Fig. 6 (bottom). In order to confirm this phase transition further, we repeat the calculations for other values of $k$, the results of which are shown in Figure 7. It is to be noted that $k=4$ is a mixed phase space with considerable chaos already present in the system.
@@
Table 1 lists the critical disorder strength $w_{c}$ along with the critical exponents $\nu$ and $\zeta$ for different values of the interaction strength or chaos parameter $k$. Intuitively, the chaotic dynamics at larger $k$ values will require smaller disorder to take the dynamics to the full Hilbert space, which is also what we observe. Interestingly, we also find a weak parameter dependence on the critical exponent $\nu$ for this phase transition. The exponent $\nu$ lies within a range of 0.6 to 0.3 for the parameters studied. Such a variation seems to be a characteristic of a finite size disordered system such as the well studied many-body-localization Wang et al. (2021). We also find that in most cases, $\nu\sim\zeta$. The quality of the data collapse for different values of $k$ confirm the presence of a second order phase transition. At the same time, we do observe that identifying correct $w_{c}$ values for larger $k$ is difficult since $w_{c}$ then is very small and a single crossing point for all system sizes is difficult to obtain. On the other hand when $k\rightarrow 0$, the dynamics can get highly non-ergodic and these parameters need more careful examinations. We have further confirmed this phase transition using the entanglement entropy $\overline{S}_{N/2}$ with a similar finite size scaling given in Eq. 10 resulting to $\nu=0.67$, $\zeta=0.76$ and $w_{c}=1.8$ when $k=1$, as shown in Fig. 8. We do acknowledge the fact that these numerics are limited by finite size systems and may not fully capture the behavior in the thermodynamic limit. This may also be the reason for slightly different values of $w_{c}$ obtained using $\overline{S}_{N/2}$ and $\overline{\langle J^{2}\rangle}$.
@@
Table 1: Critical disorder strength and exponents $\nu$, $\zeta$ for different chaos parameter, $k$. We have used $\overline{\langle J^{2}\rangle}$ with $N=12,14$ and $16$ to identify the critical points with time averaging ranging from $n=10^{5}$ to $3\times 10^{5}$ for 100 different realizations.
@@
Figure 8: (a) Plot of $\overline{S}_{N/2}/N^{\zeta/\nu}$ vs $w$ for $k=1$ corresponding to different system sizes cross each other at $w=w_{c}\sim 1.8$. Inset shows the collapse of the data with $\nu=0.67$ and $\zeta=0.76$.
It is also observed that the collapse is relatively good when the clean system is either predominantly chaotic or predominantly regular. For a mixed phase space, with significant fraction of both regular and chaotic regions, say at $k=3$, the quality of the scaling seems to depend on the initial state. The value in the table 1 for $k=3$ and $k=4$ corresponds to chaotic regions. For regular regions in these cases, we could not observe scaling behaviour. This may be attributed to the small values of angular quantum number $j$ (proportional to system size), that we are forced to restrict ourselves to, as well as the complexity of mixed phase space (see for example Madhok (2015) for cases when the initial state in a kicked top is at the border between chaos and regular regions) and the influence of phenomena such as chaos assisted tunneling Tomsovic and Ullmo (1994); S.Keshavamurthy and P.Schlagheck ; Krithika et al. (2023), that warrants a detailed separate study.
III.2 Nongeneric initial states
The above calculations were performed for a generic initial state specified by a $(\theta,\phi)$ which has no special significance. It is interesting to study the effect of changing this initial state on the nonequilibrium dynamics and on the critical behavior. For example, in the solvable case we found that for all initial states in the $y-z$ plane of the Bloch sphere, $\langle J^{2}(n)\rangle$ eventually saturates to the RMT value, independent of the strength of the disorder. Figure 9 shows the scaling of the saturation value $\langle J^{2}\rangle$ for four different initial coherent states (see Fig. 2 for their location in phase space), including $(\pi/2,\pi/2)$, $(\pi/2,0)$, and two others $(\theta,\phi)=(\pi/2,1.9)$, $(\theta,\phi)=(\pi/2,-0.56)$, which are somewhat close to $(\pi/2,\pi/2)$ and $(\pi/2,0)$, respectively. We observe a shift in $w_{c}$ for certain initial states as compared to the one reported in table 1.
To understand this initial state dependence behavior better, at least qualitatively, let us start from an initial state given by $(\theta,\phi)=(\pi/2,\pi/2)$, which is an eigenstate of $\sigma_{y}$, so that the many body state from Eq. 5 is fully delocalized along the $x$-direction, in which the Hamiltonian has both the interaction and the disorder. In other words the quantum coherence, as measured by the off-diagonal matrix elements of the initial state is the largest possible in the $\sigma^{x}$ basis. Several recent studies have shown the importance of initial state coherence as a resource for the process of thermalization Styliaris et al. (2019); Pulikkottil et al. (2023); Mitra and Srivastava (2023). Thus these initial states with large coherence tend to delocalize preferentially. In the absence of disorder, the classical limit corresponding to these initial states, for $k=1$, are stable fixed points, and it is interesting that the quantum effects of disorder tend to preferentially delocalize these states. This is consistent with a lower value of the critical strength $w_{c}$ of the disorder. However, the crossing points in Fig. 9 are drifting most significantly in this case towards smaller values of $w_{c}$ as $N$ increases, and the scaling does not seem to work well. But with the system sizes we are able to go to, $w_{c}$ is not larger than $\approx 1.2$.
On the other hand when $(\theta,\phi)=(\pi/2,0)$, the initial state is an eigenstate of $\sigma_{x}$, with no off-diagonal matrix elements in $\sigma^{x}$ basis, or zero quantum coherence. In this case, stronger disorder along the $x$ interaction direction is required to equilibriate in the full Hilbert space. The scaling behavior seems to be more convincing as well, and the critical disorder strength $w_{c}$ is not larger than $\approx 2.1$. Classically this corresponds still to a stable point, however, there is a small resonance island with nearby unstable points. Thus, the coherence of the initial states seems to play a crucial role when the dynamics is regular. We have numerically verified this argument using other values of $(\theta,\phi)$ as well. The initial states with $(\theta,\phi)$ closer to the $(\pi/2,0)$ has a larger $w_{c}$ as compared to the ones closer to $(\pi/2,\pi/2)$, but with similar critical exponents, as shown in Fig. 9 c, d.
Figure 9: Finite size scaling for $k=1$ and different initial conditions $|\theta,\phi\rangle$: (a) $|\pi/2,\pi/2\rangle$ shows $w_{c}=1.2$, $\nu=0.55$ and $\zeta=0.58$, (b) $|\pi/2,0\rangle$ with $w_{c}=2.1$, $\nu=0.66$ and $\zeta=0.47$, (c) $|\pi/2,1.9\rangle$, with $w_{c}=1.1$, $\nu=0.54$ and $\zeta=0.64$. This initial condition is closer to $(\pi/2,\pi/2)$, and hence has a smaller $w_{c}$, (d) $|\pi/2,-0.56\rangle$ with $w_{c}=2.3$, $\nu=0.57$ and $\zeta=0.58$ is similar to Fig. 6, since both the initial states are closer to $(\pi/2,0)$.
IV Conclusion
@@
In summary, this work point towards the existence of a critical strength of disorder at which there is a second order phase transition from area law entangled states with total angular momentum scaling as $N$, to volume law entangled states with total angular momentum scaling as $\sqrt{N}$. The critical disorder strength shows clear dependence on the dynamics being regular or chaotic in the disorder-free permutation symmetric space. The exponents on the other hand seem to show a weaker dependence. Most of the initial states that we checked do show a good finite size scaling behavior with roughly the same exponents as mentioned in Table 1. We also understand the importance of finite sizes in these studies which may effect the results in the thermodynamic limit.
@@
From an applications perspective, our results show that for disorder less than the critical strength, permutation symmetric systems are likely to be reliably used for their intended application, being at least not volume law entangled in the full Hilbert space. It is possible that the proposed phase transition be experimentally verified by extending few qubit experimental kicked top studies Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019), wherein disorder is anyway present. In fact, a disordered all-to-all interacting system, simulating the SYK model, has been recently experimentally studied using a combination of atomic cloud in a cavity and controllable light shift in the large detuning limit Sauerwein et al. (2023).
@@
Among future directions, it will be interesting to explore if breaking permutation symmetry through other methods, say by introducing a power law interaction such as in Russomanno et al. (2021), will also exhibit a similar phase transition as induced by disorder.
Acknowledgements
@@
MC and UD acknowledge the HPC facility Chandra at IIT Palakkad where the computations were carried out. AL thanks Sylvia Pappalardi for discussions at an early stage of this work, at the program on “Dynamical Foundations of Many-Body Chaos”, Institut Pascal, Université Paris-Saclay, during March-April 2023, whose organizers are also gratefully acknowledged. Funding support from the Department of Science and Technology, Govt. of India, under Grant No. DST/ICPS/QuST/Theme-3/2019/Q69 and MPhasis for supporting the CQuiCC center are gratefully acknowledged by AL.